# access point PASS pinching-antenna system

### v1

@@ I Introduction
Following the DOCOMO demonstration in [1], pinching-antenna systems have been proposed as a new technology to realize an analog front-end with flexible-antenna arrays [2]. A PASS consists of a long dielectric waveguide in which electromagnetic (EM) waves propagate with minimal attenuation. The EM signal is radiated from this waveguide using the so-called pinching elements which are freely moving across the waveguide [3]. The long scale of the waveguide allows the elements to place close to users and thereby drastically reduce the path-loss [3, 4]. Early studies on PASS have shown that by optimally placing and activating pinching elements, one can push the system toward a near-wired behavior, where the path-loss is dominated by the the free-space propagation at the last-meter transmission [2]. Motivated by its high degrees of freedom, several recent lines of work have investigated design aspects of PASS. The work in [5, 6] studies the integrability of this technology to multiple-input multiple-output (MIMO) systems, developing low-complexity algorithms for hybrid uplink and downlink beamforming with PASS. The study in [7] characterizes sevral performance metrics under realistic waveguide attenuation and path-loss models. Enhancing physical layer security using the high flexibility of PASSs is studied in [8].
@@ I-B Contributions and Organization
In this work, we study both dynamic and static designs for energy efficient transmission in PASSs. Our key contributions are three-fold: ($i$) We model the EE of PASSs taking into account the EM propagation in these systems. We show that conventional formulation based on earlier power distribution models is not applicable to PASSs. In fact, the power allocation in these systems is implicitly controlled through tuning of the pinching couplings and locations. ($ii$) Using our EE model, we formulate energy-efficient transmission in PASSs. The design variables in this problem vary at different rates: while per-user power allocation can be tuned at symbol-level, the coupling coefficients of the elements can be tuned only statically. We hence consider two realistic design scenarios: ($i$) a dynamic scenario, in which both user power allocation and pinching locations are adjusted at the rate of network coherence, i.e., they change as users change their locations, and the coupling lengths of the elements are tuned statistically. ($ii$) A static case in which only the user power allocation is adjusted at the coherence rate, and the pinching locations and couplings are tuned for optimal ergodic EE. (iiiiii) We develop a two-tier algorithmic framework, whose inner layer updates the dynamic variables at the coherence rate, while the outer layer tunes the static parameters. Through extensive numerical experiments, we validate the our analysis. Our numerical experiments demonstrate significant throughput gain achieved by dynamic tuning of pinching locations, suggesting high efficiency of PASSs in environment with slow dynamics, e.g. indoor settings with low user mobility.
@@ II-A Power Model for Downlink PASS
$s_{k,1}=\sin\left(\kappa\ell_{1}\right)e^{-j\beta x_{1}}s_{k,0}.$ (2)
@@
$s_{k,1}^{\rm r}=\cos\left(\kappa\ell_{1}\right)e^{-j\beta x_{1}}s_{k,0},$ (3)
The signal $s^{1}_{\rm in}$ can be treated as a new input fed at location $x_{1}$, whose splitting at $x_{2}-x_{1}$ specifies the radiation on pinching element $2$. By repeating this procedure, we can conclude that the signal radiated by element $n$ in time interval $k$ is [11]
@@
$A_{n}\left(\bm{\delta}\right)=\sqrt{1-\delta_{n}^{2}}\left(\prod_{i=1}^{n-1}\delta_{i}\right)$ (5)
@@
$\bar{P}^{\rm tx}=\frac{1}{K}\sum_{k=1}^{K}P_{k}^{\rm tx}=\bar{P}^{\rm in}\sum_{n=1}^{N}A_{n}^{2}\left(\bm{\delta}\right)$ (8)
@@
As mentioned, we assume that the remaining power in the waveguide is negligible. In asymmetric architectures, one can set this power to zero by matching the last pinching element, i.e., by setting $\delta_{N}=\cos\left(\kappa\ell_{N}\right)=0$. In a symmetric architecture, i.e., $\delta_{n}=\delta$ for all $n\in\left[N\right]$, this is not feasible. However, the remaining power in this case is $\delta^{2N}\bar{P}^{\rm in}$ which can always set arbitrarily close to zero if $\ell$ is chosen such that $\delta^{2N}$ is small enough for the given $N$.
In addition to the input power, there is a circuit power consumed by the radio frequency chain deployed to radiate the modulated signal in the waveguide. This is modeled as a single term $P^{\rm cir}$, as there is a single chain per waveguide. Assuming this term to be fixed, we can model the total power consumed by this PASS as
@@ III Static and Dynamic Design for PASS
The ultimate goal of this work is to develop an algorithmic approach to optimize the EE of this system. This is a hybrid design problem in which the signal-level parameters, i.e., user power allocation, and system-level parameters, i.e., location and coupling length of pinching elements, are designed jointly for optimal EE. One can readily see that the optimal design depends on the system setting, e.g., users’ location, and hence any algorithmic approach should update the design parameters at a certain rate. While signal-level parameters can be updated at high rate, the update rate of system-level parameters can be limited due to physical restriction.
@@ III-A Dynamic Design of PASS
$\text{ s.t.}\;\;\bm{\delta}\in\left[0,1\right]^{N},\mathbf{x}\in\mathbb{L},\text{ and }\sum_{k=1}^{K}P_{k}^{\rm in}\leq KP_{0}$ ($\mathrm{C}$)
@@ III-B Static Design of PASS
$\max_{\bm{\delta},\mathbf{x}}\;$ $\mathbb{E}_{\mathbf{u}_{1},\ldots,\mathbf{u}_{K}}\left\{\max_{\mathbf{P}}\mathcal{E}\left(\mathbf{P}^{\rm in},\mathbf{x},\bm{\delta}\right)\right\}\text{ s.t.}\;\;\;(\mathrm{C})$ ($\mathcal{S}$)
@@ IV-A Power Allocation
$\max_{\mathbf{P}^{\rm in}}\frac{\sum_{k}\log\left(1+a_{k}P_{k}^{\rm in}\right)}{KP^{\rm cir}+\sum_{k}P_{k}^{\rm in}}\text{ s.t.}\;\;\sum_{k}P_{k}^{\rm in}\leq P_{0}$ ($\mathcal{M}_{1}$)
@@ IV-B PASS Tuning
Assuming that the coherence interval of the system, i.e. number of transmission frames per which the location of the elements is optimized once, contains $m$ samples of locations $\mathbf{u}_{k,1},\ldots,\mathbf{u}_{k,m}$ for user $k$, we can estimate the ergodic EE as
@@ V Numerical Experiment
We next validate the proposed designs through numerical experiments and compare it against the baseline. Unless stated otherwise, $K=6$ users are randomly and uniformly distributed within a rectangular service region of size $D_{x}=50$ m and $D_{y}=20$ m, centered at $\frac{D_{x}}{2},0,0]$. The waveguide is aligned along the $x$-axis, deployed at a height of $a=3$ m, and covers the entire service area. The parameters of the simulation setup are summarized as follows: carrier frequency is $f_{\rm{c}}=28$ GHz, effective refractive index is $i_{\rm ref}=1.4$, circuit power is set to $P^{\rm{cir}}=0$ dBm, grid search resolution is $Q=10^{4}$, minimum inter-antenna distance is set to $\Delta={\lambda}/{2}$, and noise variance a is $\sigma_{k}^{2}=-90$ dBm. In iterative algorithms, user transmit powers are initialized as $P_{k}^{\rm{in}}={P_{0}}/{K}$, the power-splitting coefficients are set to $\delta_{n}=0.5$ for $n\in[N]$, and the initial locations of pinching elements are uniformly distributed along the waveguide as $x_{n}=(n-1){D_{x}}/{N-1}$. For comparison, we evaluate the proposed coupling-length–tunable PASS against a baseline configuration where all antennas employ a fixed power-splitting coefficient of $\delta_{n}=0.5$ for $n\in[N]$.
@@
Figure 2 shows the EE against the maximum transmit power $P_{0}$. For completeness, both EE-oriented and SE-oriented designs are evaluated, where the latter uses the proposed framework to directly maximize the SE. As shown, for both dynamic and static configurations, the EE-oriented design achieves higher EE than the SE-oriented counterpart, particularly in the high-SNR regime. Furthermore, the dynamic design outperforms the static one due to its ability to adaptively optimize the pinching-antenna placement for each user location. This highlights the critical role of antenna placement in enhancing system efficiency. In addition, when compared with the baseline scheme employing a fixed coupling length, the proposed PASS design with tunable coupling length significantly improves the EE especially under dynamic design, which validates the effectiveness of adaptive coupling-length control in improving system performance. For reference, results for a conventional fixed-position antenna located at the center of the service region are also included. Although the static design avoids frequent updates of antenna positions, its performance gain over traditional fixed-antenna systems remains limited, underscoring the importance of spatial adaptability in PASS.
@@
Figure 3 plots the EE as a function of the number of antennas. The figure shows that for the proposed designs, the achievable EE increases steadily with the antenna number. In contrast, for the PASS design with tunable coupling length, the EE exhibits only marginal improvement as the number of antennas grows. This occurs because, under fixed coupling length, most of the radiated energy is emitted by the first few antennas, and thus adding more antennas contributes little to overall system efficiency. Finally, in Figure 4, we plot the EE against the side length $D_{x}$. For the dynamic design, the antenna placement is optimized individually for each user within its allocated time slot to allow the system to maintain nearly constant EE even when users are distributed over a wider area. In contrast, the static design optimizes antenna placement based on the overall user distribution, and its achievable EE decreases as $D_{x}$ increases. This degradation arises, since the average user-to-antenna distance becomes larger with a wider service region, which leads to higher path loss.

### v2

@@ I Introduction
Following the DOCOMO demonstration in [4], pinching-antenna systems have been proposed as a new technology to realize an analog front-end with flexible-antenna arrays [3]. A PASS consists of a long dielectric waveguide in which electromagnetic (EM) waves propagate with minimal attenuation. The EM signal is radiated from this waveguide using the so-called pinching elements which are freely moving across the waveguide [5]. The long scale of the waveguide allows the elements to place close to users and thereby drastically reduce the path-loss [5, 7]. Early studies on PASS have shown that by optimally placing and activating pinching elements, one can push the system toward a near-wired behavior, where the path-loss is dominated by the the free-space propagation at the last-meter transmission [3]. Motivated by its high degrees of freedom, several recent lines of work have investigated design aspects of PASS. The work in [1] and [2] studies the integrability of this technology to multiple-input multiple-output (MIMO) systems, developing low-complexity algorithms for hybrid uplink and downlink beamforming with PASS. The study in [10] characterizes sevral performance metrics under realistic waveguide attenuation and path-loss models. Enhancing physical layer security using the high flexibility of PASSs is studied in [9].
@@ I-B Contributions and Organization
In this work, we study both dynamic and static designs for energy efficient transmission in PASSs. Our key contributions are three-fold: ($i$) We model the EE of PASSs taking into account the EM propagation in these systems. We show that conventional formulation based on earlier power distribution models is not applicable to PASSs. In fact, the power allocation in these systems is implicitly controlled through tuning of the pinching couplings and locations. ($ii$) Using our EE model, we formulate energy-efficient transmission in PASSs. The design variables in this problem vary at different rates: while per-user power allocation can be tuned at symbol-level, the coupling coefficients of the elements can be tuned only statically. We hence consider two realistic design scenarios: ($i$) a dynamic scenario, in which both user power allocation and pinching locations are adjusted at the rate of network coherence, i.e., they change as users change their locations, and the coupling lengths of the elements are tuned statistically. ($ii$) A static case in which only the user power allocation is adjusted at the coherence rate, and the pinching locations and couplings are tuned for optimal ergodic EE. (iiiiii) We develop a two-tier algorithmic framework, whose inner layer updates the dynamic variables at the coherence rate, while the outer layer tunes the static parameters. Through extensive numerical experiments, we validate our analysis. Our numerical experiments depict a significant throughput gain achieved by dynamic tuning of pinching locations, suggesting high efficiency of PASSs in environment with slow dynamics, e.g. indoor settings with low user mobility.
@@ II-A Power Model for Downlink PASS
$s_{k,1}=\sin\left(\kappa\ell_{1}\right)e^{-j\beta x_{1}}s_{k,0},$ (2)
@@
$s_{k,1}^{\rm r}=\cos\left(\kappa\ell_{1}\right)e^{-j\beta x_{1}}s_{k,0}.$ (3)
The signal $s_{k,1}^{\rm r}$ can be treated as a new input fed at location $x_{1}$, whose splitting at $x_{2}-x_{1}$ specifies the radiation on pinching element $2$. By repeating this procedure, we can conclude that the signal radiated by element $n$ in time interval $k$ is [11]
@@
$A_{n}\left(\bm{\delta}\right)=\sqrt{1-\delta_{n}^{2}}\left(\prod_{i=1}^{n-1}\delta_{i}\right),$ (5)
@@
$\bar{P}^{\rm tx}=\frac{1}{K}\sum_{k=1}^{K}P_{k}^{\rm tx}=\bar{P}^{\rm in}\sum_{n=1}^{N}A_{n}^{2}\left(\bm{\delta}\right),$ (8)
@@
As mentioned, we assume that the remaining power in the waveguide is negligible. In asymmetric architectures, one can set this power to zero by matching the last pinching element, i.e., by setting $\delta_{N}=\cos\left(\kappa\ell_{N}\right)=0$. In a symmetric architecture, i.e., $\delta_{n}=\delta$ for all $n\in\left[N\right]$, this is not feasible. However, the remaining power in this case is $\delta^{2N}\bar{P}^{\rm in}$ which can always set arbitrarily close to zero if the length of the elements (specifying $\delta$) is tuned properly.
In addition to the input power, there is a circuit power consumed by the radio frequency chain deployed to radiate the modulated signal in the waveguide. This is modeled as a single term $P^{\rm cir}$, as there is a single chain connected to the waveguide. Assuming this term to be fixed, we can model the total power consumed by this PASS as
@@ III Static and Dynamic Design for PASS
The ultimate goal of this work is to develop an algorithmic approach to optimize the EE of this system. This is a hybrid design problem in which the signal-level parameters, i.e., user power allocation, and system-level parameters, i.e., location and coupling length of pinching elements, are designed jointly for optimal EE. One can readily see that the optimal design depends on the system setting, e.g., users’ location. Therefore, any algorithmic approach should update the design parameters at a certain rate. While signal-level parameters can be updated at high rate, the update rate of system-level parameters can be limited due to physical restriction.
@@ III-A Dynamic Design of PASS
$\text{ s.t.}\;\;\bm{\delta}\in\left[0,1\right]^{N},\mathbf{x}\in\mathbb{L},\text{ and }\sum_{k=1}^{K}P_{k}^{\rm in}\leq KP_{0},$ ($\mathrm{C}$)
@@ III-B Static Design of PASS
$\max_{\bm{\delta},\mathbf{x}}\;$ $\mathbb{E}_{\mathbf{u}_{1},\ldots,\mathbf{u}_{K}}\left\{\max_{\mathbf{P}}\mathcal{E}\left(\mathbf{P}^{\rm in},\mathbf{x},\bm{\delta}\right)\right\}\text{ s.t.}\;\;\;(\mathrm{C}),$ ($\mathcal{S}$)
@@ IV-A Power Allocation
$\max_{\mathbf{P}^{\rm in}}\frac{\sum_{k}\log\left(1+a_{k}P_{k}^{\rm in}\right)}{KP^{\rm cir}+\sum_{k}P_{k}^{\rm in}}\text{ s.t.}\;\;\sum_{k}P_{k}^{\rm in}\leq P_{0}.$ ($\mathcal{M}_{1}$)
@@ IV-B PASS Tuning
Assuming that the coherence interval of the system, i.e. number of transmission frames, per which the location of the elements is optimized once, contains $m$ samples of locations $\mathbf{u}_{k,1},\ldots,\mathbf{u}_{k,m}$ for user $k$, we can estimate the ergodic EE as
@@ V Numerical Experiment
We next validate the proposed designs through numerical experiments and compare it against the baseline. Unless stated otherwise, $K=6$ users are randomly and uniformly distributed within a rectangular service region of size $D_{x}=50$ m and $D_{y}=20$ m, centered at ${D_{x}}/{2},0,0]$. The waveguide is aligned along the $x$-axis, deployed at a height of $a=3$ m, and covers the entire service area. The parameters of the simulation setup are summarized as follows: carrier frequency is $f_{\rm{c}}=28$ GHz, effective refractive index is $i_{\rm ref}=1.4$, circuit power is set to $P^{\rm{cir}}=0$ dBm, grid search resolution is $Q=10^{4}$, minimum inter-antenna distance is set to $\Delta={\lambda}/{2}$, and noise variance is $\sigma_{k}^{2}=-90$ dBm. In iterative algorithms, user powers are initialized as $P_{k}^{\rm{in}}={P}/{K}$ for a total power $P$, the splitting coefficients are set to $\delta_{n}=0.5$ for $n\in[N]$, and the initial pinching locations are uniformly distributed, i.e. $x_{n}=(n-1){D_{x}}/{N-1}$. For comparison, we evaluate the proposed coupling-length–tunable PASS against a baseline configuration where all antennas employ a fixed power-splitting coefficient of $\delta_{n}=0.5$ for $n\in[N]$.
@@
Figure 2 shows the EE against the maximum transmit power $P$. For completeness, both EE-oriented and SE-oriented designs are evaluated, where the latter uses the proposed framework to directly maximize the SE. As shown, for both dynamic and static configurations, the EE-oriented design achieves higher EE than the SE-oriented counterpart, particularly in the high-SNR regime. Furthermore, the dynamic design outperforms the static one due to its ability to adaptively optimize the pinching-antenna placement for each user location. This highlights the critical role of antenna placement in enhancing system efficiency. In addition, when compared with the baseline scheme employing a fixed coupling length, the proposed PASS design with tunable coupling length significantly improves the EE especially under dynamic design, which validates the effectiveness of adaptive coupling-length control in improving system performance. For reference, the results for a conventional fixed-position antenna located at the center of the service region are also included. Although the static design avoids frequent updates of antenna positions, its performance gain over traditional fixed-antenna systems remains limited, underscoring the importance of spatial adaptability in PASS.
@@
Figure 3 plots the EE as a function of the number of antennas. The figure shows that for the proposed designs, the achievable EE increases steadily with the antenna number. In contrast, for the PASS design with tunable coupling length, the EE exhibits only marginal improvement as the number of antennas grows. This occurs because, under fixed coupling length, most of the radiated energy is emitted by the first few antennas, and thus adding more antennas contributes little to overall system efficiency. Finally, in Figure 4, we plot the EE against the side length $D_{x}$. For the dynamic design, the antenna placement is optimized individually for each user within its allocated time slot. This allows the system to maintain nearly constant EE even when users are distributed over a wider area. In contrast, the static design optimizes antenna placement based on the overall user distribution, and its achievable EE decreases as $D_{x}$ increases. This degradation arises, since the average user-to-antenna distance becomes larger with a wider service region, which leads to higher path loss.